# Sparse Ellipsoidal Radial Basis Function Network for Point Cloud Surface Representation

### v2

@@ 4.2 Evaluation Metric
The evaluation metrics, including Hausdorff distance (HD)[43], Chamfer distance (CD) [13], and cosine similarity (CS) [3], are used to quantify the geometric discrepancy between the reconstructed surface (extracted from the approximated SDF) and the ground truth surface (which can be regarded as the zero-level set of the real SDF). The evaluation of reconstruction quality essentially reflects the accuracy of the SDF approximation. Suppose that $S_{t}$ and $S_{r}$ are the ground truth surface and the reconstructed surface, respectively. Both HD and CD are computed based on point-to-surface distances, where the distance from a point to a surface is defined as the minimum Euclidean distance to any triangle face on the surface. Formally, the distance from a point $x$ to surface $S$ is defined as:
$d(x,S)=\min_{f\in S}\mathrm{dist}(x,f),$ (27)
where $f$ denotes a triangular face on $S$, and $\mathrm{dist}(x,f)$ is the Euclidean distance from point $x$ to the face $f$.
The HD measures the worst-case deviation between the ground truth and the reconstructed surface, and is defined as:
@@
The Chamfer Distance (CD) measures the average deviation between the ground truth and the reconstructed surface, and is computed as:
@@
where ||$\cdot$|| ⋅ | denotes the number of uniformly sampled points from the respective surfaces.
In addition to geometric distances, normal consistency is another critical factor in assessing surface reconstruction quality. To evaluate the similarity between normal vectors, we compute the cosine similarity between the normals of corresponding nearest-neighbor points:
@@
where $v(x)$ is the surface normal at point $x$, $\text{closest}_{S_{r}}(x)$ is the point in $S_{r}$ closest to $x$.
@@
For all metrics, we uniformly sample $10^{5}$ points from each surface for evaluation.
4.3 Surface Reconstruction Results
In this experiment, we sampled 40K points for each mesh using Trimesh [67] API. We primarily utilize the Kaolin API [60] to compute the values of SDF on octree grid points for the watertight surface in SE-RBFNet, while for non-watertight surfaces, we use the point-cloud-utils API [61]. Since the SparseRBF method generates extraneous surfaces in non-surface regions on some point clouds, which affects the clarity of the visual presentation, we manually remove these redundant surfaces during visualization. However, for quantitative evaluation, we still use the original reconstructed surface to compute the evaluation metrics.
(a) Error Map and Detail Comparison of House
(b) Error Map and Detail Comparison of Dragon
(c) Error Map and Detail Comparison of Hand
(d) Error Map and Detail Comparison of xyzrgb-statuette
Figure 3: Comparison of the reconstructed surfaces of our method and the SparseRBF method on uniformly sampled point clouds. (b), (c), and (d) are from the Famous dataset, while (a) is from the Thingi10k dataset.
Figure 3 presents the comparison of reconstruction results on four point clouds: House, Dragon, Hand, and xyzrgb-statuette. These point clouds pose significant challenges due to their intricate surface details and dense data distribution. As shown in the figure, our method preserves more geometric details and achieves a surface closer to the ground truth compared to SparseRBF. Specifically, Figure 3a illustrates the reconstruction error and a detailed comparison of the House model. Our method achieves better reconstruction quality, especially in fine structures and thin-edge regions, where SparseRBF exhibits varying degrees of defects and missing features. For the Hand model, as shown in Figure 3c, the SparseRBF approach produces noticeable holes, resulting in incomplete reconstructions. Similarly, the Dragon and xyzrgb-statuette error maps further demonstrate our method’s ability to closely approximate the point cloud surface. For instance, in the xyzrgb-statuette case, the surface reconstructed using our method is much closer to the ground truth surface, whereas SparseRBF either fill small holes or break apart, leading to inaccurate reconstructions. These qualitative results validate the effectiveness of our sparse optimization strategy in accurately modeling fine-scale geometric features.
Table 1: Reconstruction Accuracy, Computation Time (in seconds) , and Sparsity Comparison Between Our Method and SparseRBF for the Point Cloud in Figure 3.
Point Cloud Method Reconstruction Metrics Sparsity Comparison
HD CD CS Time Init Basis Opt Basis
House Ours 0.3321 0.00036 0.9826 228.2 9050 4450
SparseRBF 0.5471 0.0106 0.9489 541.4 16306 4768
Dragon Ours 0.0216 0.00014 0.9928 163.9 4602 3615
SparseRBF 0.5730 0.0081 0.9346 620.3 18454 4341
Hand Ours 0.0028 0.00013 0.9916 105.2 2779 1942
SparseRBF 0.0101 0.00073 0.9752 546.0 15934 3283
xyzrgb-statuette Ours 0.0045 0.00036 0.9034 315.2 4476 3817
SparseRBF 0.6493 0.0334 0.6158 700.6 18273 7100
In Table 1, we present a comprehensive quantitative study of the reconstruction results of the different point clouds shown in Figure 3. The results show that our method achieves the better performance across all metrics, including HD, CD, and CS. These results indicate that our approach produces high-quality reconstructions with well-preserved surface details. This performance advantage is attributed to the powerful approximation capability of ERBFs, which provide a flexible and accurate surface representation. In contrast, SparseRBF shows significantly inferior performance, which can be mainly ascribed to two factors: (1) the reconstructed surface is incomplete (as shown in Figure 3d), and (2) redundant surfaces are generated in non-surface regions (as shown in Figure 3a). These problems significantly degrade the accuracy and quality of reconstruction.
In the sparsity comparison, both SparseRBF and our proposed method use an implicit function representation based on RBFs, allowing point clouds to be represented with significantly fewer basis functions. This sparsity is crucial for reducing storage and computational overhead, especially when processing large point clouds. The right side of Table 1 shows a comparison of the effective basis functions. Here, Init Basis denotes the initial basis functions. For our method, Init Basis includes both the initial and the newly added basis functions during optimization, while Opt Basis denotes the effective basis functions remaining after the optimization. Compared to SparseRBF, our method significantly reduces the number of optimized basis functions. For example, in the Hand point cloud, our method reduces the number of basis functions by approximately 40%, while for the xyzrgb-statuette model, the reduction approaches 50%. This indicates that our method achieves a higher compression ratio while maintaining superior reconstruction quality. Furthermore, the computation time of our method is about three times faster than that of SparseRBF, highlighting the effectiveness of our sparse optimization strategy.
In summary, our method demonstrates robust generalization and versatility over both simple and complex point cloud structures, while achieving superior reconstruction quality and reduced storage requirements. These results highlight the potential of our approach for a wide range of practical applications where efficiency and accuracy are paramount.
To further validate the generality and robustness of our proposed method, we performed a comprehensive evaluation on four datasets: ABC, Famous, MeshSeg, and Thingi10k.
(1) Reconstruction Accuracy.
Table 2 summarizes the average reconstruction accuracy of our method and SparseRBF on four benchmark datasets: ABC, Famous, MeshSeg, and Thingi10k. As shown in the table, our method consistently achieves significantly lower HD and CD values and higher CS scores across all datasets, indicating better reconstruction accuracy and higher surface fidelity. These results highlight the superior approximation capability of SE-RBFNet and confirm the effectiveness of our sparse ellipsoidal representation in preserving geometric details while maintaining a compact and efficient representation across a wide variety of point cloud distributions.
Table 2: Average Reconstruction Accuracy of the Point Clouds on the Four Datasets
Method Datasets ABC Famous MeshSeg Thingi10k
HD CD CS HD CD CS HD CD CS HD CD CS
Ours 0.0121 0.00029 0.8636 0.0099 0.00023 0.9710 0.0089 0.00021 0.9895 0.0132 0.00020 0.9697
SparseRBF 0.6177 0.05980 0.7790 0.3535 0.03265 0.8240 0.9825 0.10054 0.9139 0.4353 0.02990 0.8832
(2) Sparsity.
Figure 4 and Table 3 compare the average number of initial and optimized basis functions used by our method and SparseRBF across the four benchmark datasets. The ”Ratio” column indicates the proportion of optimized effective basis functions in our method compared to those in SparseRBF. The results demonstrate that our method significantly reduces the number of basis functions while maintaining superior reconstruction performance. In particular, our method requires only about 50% of the basis functions used by SparseRBF on average. This reduction not only reduces memory consumption but also accelerates computation, making our method highly efficient and practical for large-scale applications.
Figure 4: The Average Initial and Optimized Basis Functions on the Four Datasets.
Table 3: Comparison of Initial and Optimized Basis Functions for Different Datasets.
Dataset Method Init Opt Ratio
ABC SparseRBF 16880 4226 0.47
Ours 4943 2007
Famous SparseRBF 17595 4812 0.53
Ours 5633 2589
MeshSeg SparseRBF 17019 2749 0.41
Ours 3470 1148
Thingi10k SparseRBF 17060 3726 0.55
Ours 5079 2073
(3) Computational Efficiency.
The computational efficiency comparison is summarized in Table 4. Our method consistently outperforms SparseRBF in reconstruction speed across all four datasets. Specifically, our approach achieves an approximately 3–4$\times$ reduction in computation time, highlights the effectiveness of the proposed sparse optimization strategy in accelerating the reconstruction process. Combined with the accuracy results presented earlier, these findings demonstrate that our method offers a favorable trade-off between reconstruction quality, computational efficiency, and storage cost. This makes it especially suitable for real-world applications where both performance and efficiency are essential.
Table 4: Average Reconstruction Time (in seconds) of the Point Cloud on the Four Datasets
Method ABC Famous MeshSeg Thingi10k
Ours 129.66 185.72 116.28 149.42
SparseRBF 512.44 596.28 498.27 520.15
4.4 Real-world Scanned Data
(a) statue_ps_outliers Reconstruction Surface and Detail Comparison
(b) Error Colormaps and Detail Comparison of torch_ps_outliers
@@ 4.4 Real-world Scanned Data
Figure 5: Comparison of the reconstructed surfaces of our method and the SparseRBF method on two real-world scanned datasets. The point clouds are all provided by Erler et al. [13].
In the previous sections, we have analyzed the surface reconstruction results on several widely used benchmark datasets, demonstrating that our method performs well in terms of both accuracy and robustness. To further evaluate the generalizability and practical applicability of our approach, it is crucial to assess its performance on real-world scanned data.
Real-world scanned point clouds typically suffer from noise, incomplete regions, and irregular sampling densities, which pose significant challenges to surface reconstruction algorithms. To assess the robustness of our method under such conditions, we applied it to several real-world scanned datasets, including the statue_ps_outliers and torch_ps_outliers models provided by Erler et al. [13]. The corresponding reconstruction results are shown in Figure 5. As observed, our method continues to achieve better reconstruction quality compared to SparseRBF. In the detailed views shown in the second rows of Figure 5a and Figure 5b, our method achieves visually optimal results, effectively reconstructing small holes and thin structural connections, while SparseRBF tends to lose some fine surface details due to its smoother surface reconstruction. This highlights the robustness of our method in maintaining high reconstruction quality when dealing with complex real-world data.
@@
Table 5: Reconstruction Accuracy, Computation Time (in seconds) , and Sparsity Comparison Between Our Method and SparseRBF for the Point Cloud in Figure 5.
Point Cloud Method Reconstruction Metrics Sparsity Comparison
HD CD CS Time Init Basis Opt Basis
statue_ps_outliers Ours 0.0121 0.00019 0.9835 172.2 4773 3665
SparseRBF 0.6958 0.0023 0.9569 612.3 17539 3463
torch_ps_outliers Ours 0.0087 0.00021 0.9785 178.0 6342 4856
SparseRBF 0.4229 0.0018 0.9333 675.2 17851 3441
Table 5 provides a detailed quantitative evaluation of the results in Figure 5, further confirming the superior accuracy and robustness of our method. In the sparsity comparison, our method demonstrates a slightly higher number of effective basis functions compared to SparseRBF, especially in the torch_ps_outliers. This is primarily because our method adaptively adds basis functions to minimize approximation error, particularly in regions with complex structures. In contrast, although SparseRBF uses a slightly smaller number of basis functions, the reconstruction errors are much larger. Real-world point clouds often exhibit more complex geometric details (such as numerous fine wrinkles, small bumps, and subtle irregularities on the surface, as shown in Figure 5b), requiring a larger number of basis functions for accurate representation. Consequently, our method achieves remarkable robustness and reconstruction quality across various challenging scenarios.
4.5 Surface Reconstruction with Challenging Cases
This section focuses on the research of SE-RBFNet in special point cloud scenarios. The primary cases of interest include models with holes, slender and thin structures, multiple disjoint objects, and sharp shapes. Notably, SE-RBFNet handles these cases without requiring any parameter adjustments, demonstrating its robustness and strong generalization across diverse scenarios.
Figure 6 presents the corresponding reconstruction results on several challenging point clouds. In Figure 6a, a thin planar surface with a hole is reconstructed. Our method accurately preserves the hole, while SparseRBF tends to partially or completely fill it, resulting in unrealistic surfaces and the loss of critical geometric details. Figure 6b shows a thin-walled point cloud with slender features. SE-RBFNet achieves high-fidelity reconstruction, maintaining the integrity of both the thin walls and the slender components. In contrast, SparseRBF introduces numerous holes and artifacts, severely degrading the reconstruction quality. In the case of multiple disjoint objects, shown in Figure 6c, our method produces more complete and coherent surfaces, avoiding the spurious and fragmented structures observed in the SparseRBF results. This highlights the robustness of SE-RBFNet in handling disjoint geometries. Finally, Figure 6d shows a complex point cloud featuring sharp edges, slender components, and disjoint structures. Our method successfully reconstructs the shape with high geometric accuracy, preserving both fine-scale and global features, whereas SparseRBF fails to recover several sharp regions and produces overly smoothed surfaces, resulting in the loss of important geometric details. In summary, these experiments demonstrate that SE-RBFNet consistently delivers high-quality reconstructions across a variety of challenging scenarios, confirming its robustness and adaptability.
(a) Perforated Thin Sheets
(b) Thin Walls and Slender Structures
(c) Multiple Disjoint Object
(d) Slender Features and Multiple Disjoint Object
@@ 4.5 Surface Reconstruction with Challenging Cases
Figure 6: Comparison of the reconstructed surfaces of our method and the SparseRBF on challenging point clouds. (a), (b), and (c) are from the ABC dataset, while (d) is from the Thingi10k dataset.

### v4

@@ 4.2 Evaluation Metric
In our setting, SE-RBFNet aims to approximate a precomputed ground-truth SDF. To ensure a fair and meaningful evaluation, the ground-truth SDF is computed directly from the original triangle mesh geometry, avoiding intermediate errors of SDF values. It is important to note that our loss function assigns greater emphasis to points close to the surface (see Eq. 10). Therefore, accurately approximating SDF values on all grid points is not the purpose of SE-RBFNet. Instead, SE-RBFNet aims to preserve the geometry of the zero-level set (i.e., the implicit surface). Accordingly, our evaluation focuses on the geometric deviation between the zero-level set surface from the outputs of SE-RBFNet and the isosurface from the ground-truth SDF, as measured by surface-to-surface distances and normal consistency. Specifically, the network-predicted normalized SDF values are first transformed back to the physical SDF range according to Eq. (6). Using the same octree structure, the iso-surface at zero level is then extracted from the transformed ERBF-approximated SDF and compared with the ground truth surface directly obtained from the ground-truth SDF, which represents the true zero-level set. The explicit surface extraction from the SDF values on grids is performed using the method provided in [31]. We use the following metrics—Hausdorff Distance (HD) [31], Chamfer Distance (CD) [15], and Cosine Similarity (CS) [45]—to compare the zero-level set surface from the outputs of SE-RBFNet with the isosurface of the ground-truth SDF. Smaller HD/CD values and higher CS values indicate that the surfaces extracted from the approximated SDF closely match the true zero-level set, thus reflecting higher fidelity of the SDF representation of the implicit surface.
Suppose that $S_{t}$ and $S_{r}$ are the ground truth surface and the surface extracted from the approximated SDF, respectively. Both HD and CD are computed using point-to-point distances, where the distance from a point to a surface is defined as the minimum Euclidean distance to any point on that surface. Formally, the distance from a point $x$ to a surface $S$ is defined as:
$d(x,S)=\min_{y\in S}\|x-y\|_{2},$ (27)
where $y\in S$ denotes a surface point.
The HD measures the worst-case deviation between the surface $S_{t}$ and $S_{r}$, and is defined as:
@@
The Chamfer Distance (CD) measures the bidirectional average deviation between the surface $S_{t}$ and $S_{r}$, and is computed as:
@@
where $|\cdot|$ denotes the number of points sampled uniformly on the surface.
In addition to geometric distances, normal consistency is another critical factor in assessing surface representation quality. To evaluate the similarity between normal vectors, we compute the cosine similarity between the normals of corresponding nearest-neighbor points:
@@
where $v(x)$ is the surface normal at point $x$, $\text{closest}_{S_{r}}(x)$ is the point in $S_{r}$ closest to $x$. For all the above metrics, $10^{5}$ points are uniformly sampled from each surface for evaluation.
4.3 Sparsity and Parameter Efficiency Evaluation
(a) Error Map + Details (Horse)
(b) Error Map + Details (Hand)
(c) Error Map + Details (Person)
(d) Error Map + Details (Dragon)
(e) Error Map + Details (statue_ps_outliers)
(f) Error Map + Details (torch_ps_outliers)
@@ 4.3 Sparsity and Parameter Efficiency Evaluation
Figure 3: Comparison of implicit surface approximation results obtained by our method and SparseRBF. (a), (b), and (d) are from the Famous dataset; (c) is from the Thingi10k dataset; (e) and (f) correspond to two real-world scanned datasets provided by Erler et al. [15].
In this experiment, we focus on evaluating the sparsity and parameter efficiency of our method in comparison with SparseRBF [30], which also employs sparse RBFs for surface representation. For each mesh, 40K points are first sampled using the sample_surface_even function in the Trimesh [11] API, which performs area-weighted uniform sampling over the mesh surface. These sampled points are then used to construct an octree, and the ground-truth SDF values at the octree grid points (defined by the directional point-to-mesh distances) are precomputed directly from the mesh geometry using the Kaolin API [22], serving as the training samples for our network.
Figure 3 shows the extracted surfaces for several representative cases. As shown in the figure, the surfaces obtained from SE-RBFNet’s approximated SDF more closely match the ground-truth surface, effectively preserving geometric details while avoiding artifacts. Specifically, Figure 3a presents the error map and detailed comparison of the Horse model. Our method preserves more detailed features, such as the horse’s head and leg contours, whereas SparseRBF appears smoother but loses geometric details. For the Hand model (Figure 3b), the SparseRBF approach produces noticeable holes, resulting in incomplete surface representations. Similarly, the Person and Dragon cases further highlight the robustness of our approach in accurately approximating SDFs of the ground-truth implicit surface, with the Dragon model showing substantially reduced errors compared with SparseRBF. In addition, Figures 3e and 3f demonstrate results on two real-world scanned datasets. As illustrated by the local views in the second row of these figures, SE-RBFNet achieves visually optimal results, effectively preserving fine holes and thin structural connections, while SparseRBF tends to oversmooth these delicate features. These qualitative results confirm the effectiveness of our approach in accurately approximating the SDF, thereby enabling the extraction of high-fidelity implicit surfaces.
@@
Table 1: Accuracy, Time (in seconds) , and Sparsity Comparison Between Our Method and SparseRBF for Implicit Surfaces Shown in Figure 3.
Implicit Surface Method Surface Geometric Metrics Sparsity Comparison
HD CD CS Time Init Basis Opt Basis Param
Horse Ours 0.0042 0.0016 0.9958 106.9 3223 612 6120
SparseRBF 0.0081 0.0017 0.9919 809.7 17789 3490 17450
Hand Ours 0.0033 0.0015 0.9900 90.5 2683 1149 11490
SparseRBF 0.0126 0.0017 0.9773 654.2 15935 4970 24850
Person Ours 0.0044 0.0016 0.9949 132.0 6314 893 8930
SparseRBF 0.0062 0.0018 0.9886 1030.2 18127 5256 26280
Dragon Ours 0.0053 0.0016 0.9914 138.3 5076 1859 18590
SparseRBF 0.0206 0.0019 0.9769 1117.3 18455 6426 32130
statue_ps_outliers Ours 0.0092 0.0015 0.9861 136.7 5217 1192 11920
SparseRBF 0.0256 0.0018 0.9700 883.2 17540 5574 27870
torch_ps_outliers Ours 0.0082 0.0016 0.9830 143.1 6419 1386 13860
SparseRBF 0.0506 0.0020 0.9637 934.8 17852 5105 25525
In Table 1, we present a comprehensive quantitative study of the experimental results of the implicit surfaces shown in Figure 3. The results show that our method achieves better performance across all metrics, including HD, CD, and CS. The superior performance of SE-RBFNet highlights its ability to accurately represent the SDF using a sparse set of ERBFs.
Using as few basis functions as possible to represent the SDFs of implicit surfaces is crucial for reducing storage and computational overhead. This is particularly important when processing SDFs of complex surfaces with rich geometric details. As demonstrated in [30], SparseRBF achieves substantially better sparsity than prior RBF-based methods [5, 41, 48] by combining medial-axis-based center selection with a linear $\mathcal{L}_{1}$ sparsity optimization strategy. The right side of Table 1 shows a comparison of the number of effective basis functions obtained from SparseRBF and our method. Here, Init Basis denotes the number of initial basis functions. For our method, Init Basis includes both the initial and the newly added basis functions during optimization, while Opt Basis denotes the number of effective basis functions remaining after the optimization. In addition, Param denotes the number of parameters required to represent an implicit surface. Specifically, SE-RBFNet represents the surface using a sparse set of anisotropic ERBFs. Each ERBF is parameterized by its center, rotation angles, axis lengths, and a weight coefficient, totaling 10 parameters per ellipsoid. In contrast, SparseRBF approximates surfaces using a set of isotropic Gaussian RBFs, where each basis function is defined by its center coordinates, a scalar radius, and a weight, totaling $5$ parameters per RBF. Compared with SparseRBF, our method significantly reduces the number of basis functions, requiring on average only about 45% of its parameters while achieving higher accuracy in implicit surface representation. This demonstrates that SE-RBFNet achieves a higher compression ratio while maintaining superior approximation accuracy. Moreover, the computation speed of our method is on average approximately seven times faster than that of SparseRBF, fully validating the effectiveness of our sparse optimization strategy. Note that the time values reported in Table 1 refer only to the optimization/training process and do not include the time required for data preparation, extracting the explicit surface from the SDF, etc. In summary, these results show that SE-RBFNet can approximate precomputed SDFs with high fidelity while using fewer parameters and requiring less computation time, making it suitable for efficient and accurate implicit surface representation in practical applications.
Table 2: Average Geometric Approximation Accuracy, Training Time (in seconds), and Sparsity of Implicit Surface Representations on Four Datasets.
Dataset Method Surface Geometric Metrics Sparsity Comparison
HD CD CS Time Init Basis Opt Basis Param Ratio
ABC Ours 0.0071 0.0021 0.9910 119.0 4806 1170 11700 0.38
SparseRBF 0.0529 0.0049 0.9616 923.8 16920 6079 30395
Famous Ours 0.0058 0.0019 0.9846 167.8 6361 2050 20500 0.68
SparseRBF 0.0384 0.0022 0.9643 918.6 17596 5969 29845
MeshSeg Ours 0.0086 0.0024 0.9957 103.7 3619 582 5820 0.23
SparseRBF 0.0195 0.0026 0.9910 797.5 16980 4961 24805
Thingi10k Ours 0.0066 0.0018 0.9940 140.8 $60$ 1097 10970 0.43
SparseRBF 0.0191 0.0021 0.9844 895.5 17009 5086 25430
[…]
Figure 4: Comparison of average number of basis functions and parameters on four benchmark datasets.
4.4 General SDF Sparse Representation
As a general sparse implicit surface representation method, SE-RBFNet can be trained on any given SDF sample sets and directly implements sparse representation from SDF point values, regardless of their source (e.g., point clouds, triangle meshes, analytical SDFs, or existing neural SDFs). To further validate the method’s universality, we apply it to SDF data generated by the neural implicit method Neural-Singular-Hessian (NSH) [58]. Specifically, NSH is first used to predict SDF values at octree nodes of the input sampling points, upon which SE-RBFNet constructs its sparse representation. In addition to comparing with SparseRBF, we further compare against the fast RBF interpolation algorithm RBF-QNN provided by the ALGLIB library [51]. RBF-QNN adopts an automatic radius selection mechanism, keeps the centers and shapes of RBFs fixed during computation, and achieves computational complexity close to $O(N\log N)$, where $N$ denotes the number of training points defined previously. For fairness, all three algorithms are evaluated using the same input data to assess approximation quality.
(a) Surface Represented by the Approximate SDF and Its Error Map (3DBenchy)
(b) Surface Represented by the Approximate SDF and Its Error Map (tortuga)
(c) Surface Represented by the Approximate SDF and Its Error Map (00010218)
(d) Surface Represented by the Approximate SDF and Its Error Map (331105)
@@ 4.4 General SDF Sparse Representation
Figure 5: Comparison of the implicit surfaces extracted from the approximated SDFs by our method, SparseRBF, and RBF-QNN. (a) and (b) are from the MeshSeg dataset, (c) is from the ABC dataset, while (d) is from the Thingi10k dataset.
Experimental results are presented in Figure 5. For consistency, explicit surfaces are extracted under the same octree grid using the program provided in [31], and comparisons are made against the target surfaces generated from NSH under identical conditions. As shown in Figure 5, both our method and RBF-QNN closely approximate the SDF values predicted by NSH, while SparseRBF exhibits the most pronounced error map, primarily due to oversmoothing of fine surface details that results in significant deviations. Quantitative results are summarized in Table 3. Our method consistently outperforms SparseRBF in geometric accuracy metrics and achieves results on par with RBF-QNN. In terms of sparsity, our approach requires the fewest parameters, delivering superior results with an average of only about $35\%$ of the parameters used by SparseRBF. By contrast, RBF-QNN lacks sparsity, with a parameter count that scales linearly with the number of input points, resulting in the largest parameter size among the three methods. Regarding computational efficiency, our method exhibits the shortest training times on average: it is approximately $8\times$ faster than SparseRBF and even faster than RBF-QNN, while preserving high surface fidelity. Moreover, NSH itself requires average $264\,\mathrm{K}$ parameters to present the SDF on grids of the test cases in Table 3, whereas our method achieves high-fidelity sparse representation using fewer than $2,000$ basis functions, corresponding to no more than $20\,\mathrm{K}$ parameters. This indicates that, compared with NSH, our method maintains high approximation accuracy while achieving over $10\times$ parameter compression for the test cases in Table 3, greatly improving storage and transmission efficiency. More importantly, this experiment not only validates the superiority of the proposed method over traditional RBF interpolation baselines but also demonstrates its strong compatibility with advanced neural implicit representation models. In other words, SE-RBFNet provides a lightweight and efficient sparse representation for neural SDFs, effectively reducing computational and storage costs without sacrificing geometric fidelity. This highlights its potential for efficient geometric representation and resource-constrained scenarios.
Table 3: Accuracy, Training Time (in seconds), and Parameter Comparison of Our Method, SparseRBF, and RBF-QNN for Implicit Surfaces Shown in Figure 5. The ground-truth SDFs used in this evaluation are obtained from NSH [58].
Implicit Surface Method Surface Geometric Metrics Parameter Comparison
HD CD CS Time Basis Count Param
3DBenchy Ours 0.0068 0.0022 0.9891 130.7 2601 26010
SparseRBF 0.0912 0.0025 0.9725 901.6 11143 55715
RBF-QNN 0.0067 0.0021 0.9890 150.9 632023 3160116
tortuga Ours 0.0662 0.0023 0.9888 119.1 1688 16880
SparseRBF 0.0650 0.0024 0.9861 1188.4 8926 44630
RBF-QNN 0.0647 0.0022 0.9885 192.5 738927 3694635
00010218 Ours 0.0069 0.0021 0.9947 120.1 1638 16380
SparseRBF 0.2128 0.0084 0.9767 1007.9 10485 52425
RBF-QNN 0.0066 0.0021 0.9947 168.4 659887 3299435
331105 Ours 0.0065 0.0019 0.9962 130.1 1494 14940
SparseRBF 0.0109 0.0020 0.9918 1163.1 11314 56570
RBF-QNN 0.0065 0.0018 0.9963 183.1 721626 3608130